# SeCoKD: Aligning Large Language Models for In-Context Learning with Fewer Shots

### v1

@@ 3 SeCoKD Overview
The primary training objective of SeCoKD is to have the student model emulate the teacher model, which is activated by a handful of demonstrations. Concretely, let $\boldsymbol{D}=\left\{d_{1},d_{2},d_{3}...d_{n}\right\}$ denotes a set of demonstrations and $\boldsymbol{d}\subseteq\boldsymbol{D}$ denotes a subset. $\left(x,y,\theta\right)$ are the input query, true label, and model parameters, respectively. In the setting of few-shot learning we have $P_{\mathcal{M}}=\left(y\mid x,\boldsymbol{D},\theta_{\mathcal{M}}\right)$ for the model $\mathcal{M}$. After applying our training method, we showcase that the updated Model ${\mathcal{M}}^{\prime}$ also performs well with a high $P_{{\mathcal{M}}^{\prime}}=\left(y\mid x,\boldsymbol{d},\theta_{{\mathcal{M}}^{\prime}}\right)$. Since we focus on a self-distillation manner and we fine-tune the model with LoRA, the expression can be rewritten as $P_{{\mathcal{M}}^{\prime}}=\left(y\mid x,\boldsymbol{d},\theta_{LoRA},\theta_{\mathcal{M}}\right)$. As depicted in Figure 1, the whole pipeline can be divided into two steps. First, the teacher model is prompted with a set of demonstrations and a query. Each demonstration contains a question, a rationale, and an answer. The reasons to include some reasoning steps are two-fold: a) It is shown that Chain-of-Thought (CoT) prompting increases the reasoning ability of LLMs and thus the performance will be better (Wei et al., 2022; Shao et al., 2023). b) We need more tokens generated from the teacher model as supervision of the student model. For each task, we use a carefully curated demonstration set as gold samples. We then extract the reasoning part and the answer from the teacher model’s output and save them for later use.
In the second step, we randomly sample a subset of the available demonstrations, concatenate it with the same query as in the first step, and use this sequence as input for the student model. Then we apply Sequential-Level KD (Kim and Rush, 2016) to fine-tune the student model.
@@ 4 Experimental Settings
We aim to evaluate the performance of SeCoKD compared to directly supervised fine-tuning and the base model. Inspired by Wei et al. (2022), we choose 6 popular benchmarks, covering topics of arithmetic reasoning, commonsense reasoning, and symbolic reasoning. We conducted experiments with some of the most advancing LLMs. However, we could only test the models with less than 10 billion parameters due to the computation limits.
@@ 4.1 LLMs
We evaluate our method on three GPT-like auto-regressive transformer language models. We use the 4-bit quantized version to save computation resources (Dettmers et al., 2022). The Llama 2-7B (Touvron et al., 2023) is one of the most popular open-source LLMs. Llama 3-8B (AI@Meta, 2024) is the latest member in the Llama family and appears to be the SOTA in various benchmarks. We also use the Mistral-7B (Jiang et al., 2023) which leverages the sliding window attention (SWA) mechanism to handle variants sequence lengths effectively. We conducted all experiments on a single NVIDIA V100 (40G) GPU. For the training, we use the same LoRA [^1] https://huggingface.co/docs/diffusers/en/training/lora (Hu et al., 2021) configuration for all models, and the trainable parameters thus reduce to around 0.18% of the full size. All results reported are the average of three runs and training-related hyperparameters are listed in A.2.
@@ 4.2 Datasets
Dataset Rationale Multiple Choice
ARC-C curated
CSQA
SVAMP same as GSM8K
AQUA-RAT
GSM8K
COIN-FLIP
Table 1: We need to apply CoT prompting in our training pipeline. We reuse the existing demonstrations for GSM8K, COIN-FLIP, SVAMP and CSQA tasks from Wei et al. (2022). For ARC-C we manually curate 8 gold demonstrations. The full demonstrations are listed in A.3
Figure 2: Comparison of 4 methods with different shot numbers. The X-axis represents the number of demonstrations. The Y-axis shows the average accuracy of all six tasks. SeCoKD significantly outperforms the other two baselines in zero-shot and one-shot scenarios.
We evaluate all methods on 6 benchmarks listed in Table 1. For mathematical reasoning tasks, we selected three datasets. The GSM8K (Cobbe et al., 2021) contains 8.5K high-quality and diverse text-based grade school math problems. The SVAMP (Patel et al., 2021) applies different types of variations to the existing math problems and creates a more robust benchmark. In AQUA-RAT (Ling et al., 2017), the answers to the math problems are multiple choices. This introduces diversity into our experiments. For the commonsense reasoning tasks, we selected ARC-C (Clark et al., 2018b) which contains relatively difficult natural grade-school level questions, and the CSQA (Talmor et al., 2019) which utilized crowd-workers to create multiple-choice questions that cover a wide range of topics. We chose the Coin-Flip dataset introduced by Wei et al. (2022) for the symbolic reasoning task. In this task, the model is asked if a coin is still heads-up after $n$ people flip it. For each task, we randomly sample 800 pieces of data for training and 200 for testing.
Figure 3: Few-Shot performance on each task. The X-axis represents the number of demonstrations. Our methods SeCoKD-S and SeCoKD-M perform much better in zero-shot and one-shot compared to the two baselines.
@@ 5.1 Results for Few-Shot ICL
Figure 2 shows an overall comparison between SeCoKD and baseline methods. The underlying model here is Llama 3-8B and more results regarding different LLM structures can be found in A.1. First, while all fine-tuned models perform better than the base model, the two variants SeCoKD-S and SeCoKD-M are better than SFT in most scenarios. We observe the largest margins in the zero-shot case, meaning that context information is successfully compressed. Second, we notice that the difference between the SeCoKD-S and SeCoKD-M is quite small. This means that in the distillation process, the model does not need a strong initial context to align with the guidance from the teacher. In the following experiments, we only use the SeCoKD-S to reduce computational resources. Last, starting from Four-Shot, adding more demonstrations seems to have no more positive impact on the performance of all methods. This observation is consistent with the study from Min et al. (2021). In their work, ICL brings only marginal improvements also after around 4 demonstrations. This indicates that there is a performance upper bond for the model that can be lifted by training, not by ICL.
@@ 7 Limitations
Appendix A Example Appendix
A.1 Further Results on Few-Shot Learning

### v2

@@ 3 SeCoKD Overview
The primary training objective of SeCoKD is to have the student model emulate the teacher model, which is activated by a handful of demonstrations. Concretely, let $\boldsymbol{D}=\left\{d_{1},d_{2},d_{3}...d_{n}\right\}$ denotes a set of demonstrations and $\boldsymbol{d}\subseteq\boldsymbol{D}$ denotes a subset. $\left(x,y,\theta\right)$ are the input query, true label, and model parameters, respectively. In the setting of few-shot learning we have $P_{\mathcal{M}}=\left(y\mid x,\boldsymbol{D},\theta_{\mathcal{M}}\right)$ for the model $\mathcal{M}$. After applying our training method, we showcase that the updated Model ${\mathcal{M}}^{\prime}$ also performs well with a high $P_{{\mathcal{M}}^{\prime}}=\left(y\mid x,\boldsymbol{d},\theta_{{\mathcal{M}}^{\prime}}\right)$. Since we focus on a self-distillation manner and we fine-tune the model with LoRA, the expression can be rewritten as $P_{{\mathcal{M}}^{\prime}}=\left(y\mid x,\boldsymbol{d},\theta_{LoRA},\theta_{\mathcal{M}}\right)$. As depicted in Figure 1, the whole pipeline can be divided into two steps.
First, the teacher model is prompted with a set of demonstrations (demonstration pool) and a query. Each demonstration contains a question, a rationale, and an answer. The creation of the demonstration pool is detailed in C. The reasons to include some reasoning steps are two-fold: a) It is shown that Chain-of-Thought (CoT) prompting increases the reasoning ability of LLMs and thus the performance will be better (Wei et al., 2022; Shao et al., 2023). b) We need more tokens generated from the teacher model as supervision of the student model. For each task, we use a carefully curated demonstration set as gold samples. We then extract the reasoning part and the answer from the teacher model’s output and save them for later use.
Second, we randomly sample a subset of the available demonstrations, concatenate it with the same query as in the first step, and use this sequence as input for the student model. Then we apply Sequential-Level KD (Kim and Rush, 2016) to fine-tune the student model.
@@ 4 Experimental Settings
We aim to evaluate the performance of SeCoKD compared to directly supervised fine-tuning and the base model. Inspired by Wei et al. (2022), we choose 6 popular benchmarks, covering topics of arithmetic reasoning, commonsense reasoning, and symbolic reasoning. We conducted experiments with some of the most advancing LLMs. However, we could only test the models with less than 10 billion parameters due to the computation limits. Our code is publicly available on Github.
@@ 4.1 LLMs
We evaluate our method on three GPT-like auto-regressive transformer language models. We use the 4-bit quantized version to save computation resources (Dettmers et al., 2022). The Llama 2-7B (Touvron et al., 2023) is one of the most popular open-source LLMs. Llama 3-8B (AI@Meta, 2024) is the latest member in the Llama family and appears to be the SOTA in various benchmarks. We also use the Mistral-7B (Jiang et al., 2023) which leverages the sliding window attention (SWA) mechanism to handle variants sequence lengths effectively. We conducted all experiments on a single NVIDIA V100 (40G) GPU. For the training, we use the same LoRA [^1] https://huggingface.co/docs/diffusers/en/training/lora (Hu et al., 2021) configuration for all models, and the trainable parameters thus reduce to around 0.18% of the full size. All results reported are the average of three runs and training-related hyperparameters are listed in B.
@@ 4.2 Datasets
Figure 2: Comparison of 4 methods with different shot numbers. The X-axis represents the number of demonstrations used for inference. The Y-axis shows the average accuracy of all six tasks. SeCoKD significantly outperforms the other two baselines in zero-shot and one-shot scenarios.
We evaluate all methods on 6 datasets covering 3 types of reasoning tasks. For mathematical reasoning, we selected three datasets. The GSM8K (Cobbe et al., 2021) contains 8.5K high-quality and diverse text-based grade school math problems. The SVAMP (Patel et al., 2021) applies different types of variations to the existing math problems and creates a more robust benchmark. In AQUA-RAT (Ling et al., 2017), the answers to the math problems are multiple choices. This introduces diversity into our experiments. For the commonsense reasoning tasks, we selected ARC-C (Clark et al., 2018b) which contains relatively difficult natural grade-school level questions, and the CSQA (Talmor et al., 2019) which utilized crowd-workers to create multiple-choice questions that cover a wide range of topics. We chose the Coin-Flip dataset introduced by Wei et al. (2022) for the symbolic reasoning task. In this task, the model is asked if a coin is still heads-up after $n$ people flip it. For each task, we randomly sample 800 pieces of data for training and 200 for testing.
4.3 Training SFT model
The key of SeCoKD is to use teacher-generated rationale to align the target task with the student model. As a comparison, we train a separate model with normal SFT to exclude teacher supervision by replacing teacher-generated supervision with standalone rationales that either can be found in the dataset or created. For more details please refer to D. Specifically, we follow the step 2 of training SeCoKD with the following changes: a) For demonstration we use the whole demonstration pool D instead of a subset of it. b) We use standalone rationales and answers instead of generated ones. With such a training schema we guarantee that the SFT model uses the same amount of data as SeCoKD and thus isolates the effect of the knowledge distillation.
@@ 5 Results and Discussion
For inference and evaluation, we test the zero-shot and few-shot learning ability by using 0 to 8 demonstrations in the same demonstration pool.
Figure 3: Few-Shot performance on each task. The X-axis represents the number of demonstrations used for inference. Our methods SeCoKD-S and SeCoKD-M perform much better in zero-shot and one-shot compared to the two baselines.
@@ 5.1 Results for Few-Shot ICL
Figure 2 shows an overall comparison between SeCoKD and baseline methods. The underlying model here is Llama 3-8B and more results regarding different LLM structures can be found in A. First, while all fine-tuned models perform better than the base model, the two variants SeCoKD-S and SeCoKD-M are better than SFT in most scenarios. We observe the largest margins in the zero-shot case, meaning that context information is successfully compressed. Second, we notice that the difference between the SeCoKD-S and SeCoKD-M is quite small. This means that in the distillation process, the model does not need a strong initial context to align with the guidance from the teacher. In the following experiments, we only use the SeCoKD-S to reduce computational resources. Last, starting from Four-Shot, adding more demonstrations seems to have no more positive impact on the performance of all methods. This observation is consistent with the study from Min et al. (2021). In their work, ICL brings only marginal improvements also after around 4 demonstrations. This indicates that there is a performance upper bond for the model that can be lifted by training, not by ICL.
@@ 7 Limitations
Appendix A Further Results on Few-Shot Learning
@@ Appendix C Demonstration Pool
Appendix D Rationale for SFT training
For the datasets GSM8K and AQUA-RAT the creators provide rationales for each sample. For the rest, we prompt GPT3.5 with query and the answer to generate the corresponding rationales. The template is: "The answer to the question {query} is {answer}. Please generate a short rationale to justify the answer."